OpenStudy (anonymous):

-5y - 9 = y - 1 Possible answers: 1) -1/2 2) -2 1/2 3) -2 4) -2/5

OpenStudy (akashdeepdeb):

-5y - 9 = y - 1 -5y - y = -1 + 9 -6y = 8 y = -8/6 y = -4/3 Understood? :)

OpenStudy (anonymous):

Yup, thanks